OpenStudy (anonymous):

use the definition of derivative to find f'(x)= x/2x+11

OpenStudy (anonymous):

x/2x+11= 1/2+11=11,5 ! that's what i think!

OpenStudy (anonymous):

that wasn't helpful at all,was it?

OpenStudy (anonymous):

Are u trying to integrate? Or what, I don't understand the question...